OpenStudy (anonymous):

please help and can you explain step by step for this problem (y+1/y-2)' Also can anybody tell me if I'm right with the other problem down at the bottom please?

zepdrix (zepdrix):

This is the problem? Finding this derivative?\[\Large\bf\sf \left(\frac{y+1}{y-2}\right)'\]

OpenStudy (anonymous):

Hold on one moment please.

OpenStudy (anonymous):



OpenStudy (anonymous):

Thats the problem.

zepdrix (zepdrix):

Are both the 1 and x in the numerator? If so, you need to learn to use brackets to make it clearer XD lol

OpenStudy (anonymous):

Yes they are.

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{1+x}{\sqrt x}\right)'\]We can either quotient rule right away, or we can separate this into two fractions beforehand.

zepdrix (zepdrix):

What method would make more sense to you? :U

OpenStudy (anonymous):

Can you do it both ways?

zepdrix (zepdrix):

Here is our setup for quotient rule:\[\Large\bf\sf \left(\frac{1+x}{\sqrt x}\right)'\quad=\quad \frac{\color{royalblue}{\left[1+x\right]'}\sqrt x-(1+x)\color{royalblue}{\left[\sqrt x\right]'}}{\left[\sqrt x\right]^2}\]

zepdrix (zepdrix):

We have to take the derivative of the blue parts.

OpenStudy (anonymous):

Okay.

zepdrix (zepdrix):

Do you understand how to calculate the derivative of the first blue part? For the second part this is an important derivative to remember.\[\Large\bf\sf \frac{d}{dx}\sqrt{x}\quad=\quad \frac{1}{2\sqrt{x}}\]You can write sqrt(x) as x^(1/2) and apply the power rule if you prefer. But this one shows up so often that it's good to keep in your back pocket.

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{1+x}{\sqrt x}\right)'\quad=\quad \frac{\color{royalblue}{\left[1+x\right]'}\sqrt x-(1+x)\color{orangered}{\left[\dfrac{1}{2\sqrt x}\right]}}{x}\]

OpenStudy (anonymous):

So that's the 1st part?

zepdrix (zepdrix):

What? 0_o That's... the second part.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So my answer is in the red? or I have to divide that by x?

zepdrix (zepdrix):

I don't understand these questions you're asking... Have you learned quotient rule? :(

OpenStudy (anonymous):

I know it but I forgot it. :/

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{u}{v}\right)'\quad=\quad \frac{\color{royalblue}{u'}v-u\color{royalblue}{v'}}{v^2}\] I told you the derivative of the second term, \[\Large\bf\sf \left(\frac{u}{v}\right)'\quad=\quad \frac{\color{royalblue}{u'}v-u\color{orangered}{v'}}{v^2}\]and plugged it in. The denominator is x because we had \(\Large\bf\sf (\sqrt x)^2\) which equals x.

OpenStudy (anonymous):

So the answer is x?

zepdrix (zepdrix):

answer..?

zepdrix (zepdrix):

the denominator is x, yes..

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So that's all or there's more steps?

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{1+x}{\sqrt x}\right)'\quad=\quad \frac{\color{royalblue}{\left[1+x\right]'}\sqrt x-(1+x)\color{orangered}{\left[\dfrac{1}{2\sqrt x}\right]}}{x}\]You also need the derivative of the first blue part. Do you know how to differentiate 1+x?

OpenStudy (anonymous):

No.

zepdrix (zepdrix):

Derivative of 1?

OpenStudy (anonymous):

I'm not understanding this.

OpenStudy (anonymous):

So the derivative of 1 is -1?

zepdrix (zepdrix):

This problem seems to be way beyond your skill level :( You should start with some easier questions. No the derivative of 1 is not -1.

zepdrix (zepdrix):

The derivative of a constant is zero.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So what do I do with the remaining numbers and stuff of the problem?

zepdrix (zepdrix):

Maybe the other way would make more sense to you.\[\Large\bf\sf \frac{1+x}{\sqrt x}\quad=\quad \frac{1}{\sqrt x}+\frac{x}{\sqrt x}\]Then rewrite the roots as rational expressions,\[\Large\bf\sf =\frac{1}{x^{1/2}}+\frac{x}{x^{1/2}}\]Rules of exponents allows us to bring the first term up to the numerator by applying a negative to the exponent,\[\Large\bf\sf =x^{-1/2}+\frac{x}{x^{1/2}}\]And for the second term, when we divide things that have similar bases, we `subtract` the exponents.\[\Large\bf\sf =x^{-1/2}+x^{1-1/2}\quad=\quad x^{-1/2}+x^{1/2}\]And from there you can just apply the power rule to each term. It requires you to remember a lot of your exponent rules though, so maybe it's tougher for you.

OpenStudy (anonymous):

Can you explain how to use the power rule?

zepdrix (zepdrix):